OpenStudy (anonymous):

Algebra 2 help? I'm stumped on this very simple combining like terms problem and can't seem to get it. I'm supposed to get 5x^2/2 and can't figure out how. Problem: x + x^2/2 + 2x^2 - x.

OpenStudy (anonymous):

x^2 is over 2 or x+x^2 is over 2?

OpenStudy (anonymous):

X + (x*2/2) + 2x^2 - x.

OpenStudy (espex):

When you add your \(\frac{x^2}{2}\) and your \(2x^2\), you must give them the same denominator, giving you \(\frac{x^2}{2} + \frac{4x^2}{2}\) and your single 'x' terms cancel.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (espex):

You're welcome.